OpenStudy (anonymous):

How would I graph the following if I dont know the x or y value? 3x-5y=10

OpenStudy (bahrom7893):

it's simple: -5y = -3x + 10 y = (3/5)x - 2

OpenStudy (bahrom7893):

now suppose y = 0, what's x? 0 = (3/5)x - 2 2 = (3/5)x 10 = 3x x = 10/3 So, one point: (10/3,0)

OpenStudy (bahrom7893):

Another point.. suppose x = 0; y = 0-2 = -2 So now you have another point: (0,-2). plot the two points and connect them

OpenStudy (anonymous):

thank you